Mem. Differential Equations Math. Phys. 12(1997), 12

THE INTERNATIONAL SYMPOSIUM ON DIFFERENTIAL EQUATIONS AND MATHEMATICAL PHYSICS DEDICATED TO THE 90th BIRTHDAY ANNIVERSARY OF ACADEMICIAN I. VEKUA

The Symposium was held in Tbilisi, Georgia, in June 21–25, 1997. It was organized by the Georgian Academy of Sciences and I. Javakhishvili Tbilisi State University with the support of President of Georgia Mr. Eduard Shevardnadze.

The Scientific Committee was formed by: A. Tavkhelidze (Georgia) – Chairman, M. Atiyah (UK), B. Bojarski (Poland), R. Finn (USA), J.-L. Lions (France), G. I. Marchuk (Russia), Yu. A. Mitropol’skiĭ (Ukraine), O. A. Oleĭnik (Russia).

The members of the Organizing Committee were: I. Kiguradze (Chairman), T. Burchuladze, D. Gordeziani, G. Jaiani (Deputy Chairman), G. Kharatishvili, A. Khelashvili, V. Kokilashvili, R. Kordzadze, J. Lominadze, E. Obolashvili, N. Partsvania (Scientific Secretary), T. Shervashidze (Scientific Secretary), I. Tavkhelidze, N. Vakhania.

The Symposium had the following three sections:

• ordinary differential equations;

• partial differential equations;

• mathematical physics.

The program of the Symposium consisted of 30 minute invited talks, 15 minute short communications and poster sessions.

30 minute invited talks at the Symposium were delivered by the following participants:

a) at the section of ordinary differential equations: N. V. Azbelev (Russia), M. Drakhlin (Israel), L. Hatvani (Hungary), G. Kharatishvili (Georgia), I. Kiguradze (Georgia), T. Kusano (Japan), V. Lakshmikantham (USA), S. Leela (USA), V. P. Maksimov (Russia), F. G. Maksudov (Azerbaijan), F. Neuman (Czech Republic), B. Půža (Czech Republic), P. E. Ricci (Italy), I. P. Stavroulakis (Greece), T. Tadumadze (Georgia);

b) at the section of partial differential equations: H. Begehr (Germany), B. Bojarski (Poland), D. Cioranescu (France), R. Finn (USA), J. Gvazava (Georgia), A. G. Kartsatos (USA), S. Kharibegashvili (Georgia), V. Kokilashvili (Georgia), E. Obolashvili (Georgia), V. Paatashvili (Georgia), J.-P. Puel (France), A. Tesei (Italy);

c) at the section of mathematical physics: R. Bantsuri (Georgia), T. Burchuladze (Georgia), A. Cialdea (Italy), D. Gordeziani (Georgia), G. Jaiani (Georgia), G. Khimshiashvili (Georgia), A. I. Koshelev (Russia), M. M. Lavrent’ev (Russia), T. Meunargia (Georgia), D. Natroshvili (Georgia), T. Vashakmadze (Georgia), E. Wegert (Germany).

The invited lecturers N. A. Izobov (Belarus), L. A. Medeiros (Brazil), N. Kh. Rozov (Russia), W. Tutschke (Austria), M. Tvrdý (Czech Republic), and Z. D. Usmanov (Tajikistan) could not come to Tbilisi but sent their reports for publication.

The jubilee session was held on June 23 in Tbilisi, in Z. Paliashvili National Theatre of Opera and Ballet. It was attended not only by the participants of the Symposium but also by the broad public. President of Georgia Mr. Eduard Shevardnadze also took part in this jubilee session. Another jubilee session was held on June 26 in the town of Zugdidi, where I. Vekua passed his school years.

N. Partsvania, T. Shervashidze